# ImageSet2Text: Describing Sets of Images through Text

### v1

@@ B.1 Dataset Composition and Creation
To retrieve relevant image sets from each of the two data sources, we employed different techniques.
For the ConceptualCaptions dataset [55], we grouped images that shared the same caption and applied a filter to retain only those sets with more than 100 available links in the metadata of the original dataset. This process resulted in 287 distinct captions. After downloading the images, we found that not all links were accessible. We further filtered the sets, keeping only those with at least 50 images, which resulted in 125 unique sets. We then manually reviewed all the sets, removing 9 sets that contained either duplicate images, broken images, or images that did not correspond to the caption. Additionally, we deleted any duplicate images within the remaining 116 sets. As a result, we obtained 116 image sets, with sizes ranging from 50 to 3,342 images, for a total of 23,412 images.
The initial WikiArt dataset [57] does not include explicit reference captions on which we could group the images, but it includes metadata about the artist, style, and genre. The possible values for these attributes are listed in Tab. 5. Using this information, we generated captions for the images based on the following rules: when aggregating by genre and style, the caption format is “$<genre>$ in $<style>$ style”; when considering the artist as well, the caption format becomes “$<genre>$ by $<artist>$ in $<style>$ style”. For instance, possible captions are: “Landscapes in Romanticism Style” or “Religious paintings by Edgar Degas in Impressionism Style”. To ensure sufficient data for analysis, we filtered the groups based on the number of images: groups with only two attributes (style and genre) were kept if they contained more than 499 images, while groups with three attributes (style, genre and artist) required a minimum of 50 images. After downloading the images and removing duplicates, we retained only the groups with more than 49 images. From this process, we obtained 53,707 images distributed across 105 groups. The sizes of these groups range from 50 to 4,112 images.
@@ B.2 Baseline Methods
The baselines we compare against for this task, namely BLIP-2 [31], LLaVA-1.5 [36], and GPT-4V [1], are not designed to process multiple images simultaneously. Therefore, we conducted experiments using different settings to enable a comparison with ImageSet2Text. These settings are referred to as (a) the grid setting, (b) the group embedding setting, and (c) the summary setting. A visual summary of these settings is provided in Fig. 6.
@@
Figure 6: Settings of utilizing VQA models to generate group captions. Light colors denote images (blue), image embeddings (orange) and output text (green) on the level of individual images, dark colors denote aggregation on a group level. (a) is the grid setting where images are put into a collage to depict the group, (b) is the group embedding setting where embeddings of individual images after the vision encoder are averaged and (c) is the summary setting where an additional LLM is used to generate a group caption from the captions of individual instances.
@@
In the grid setting, we selected images from the groups, arranged them into grids, and used these grids as inputs for the baseline models. Due to resolution constraints on the input size of the respective vision encoders, we included only a subset of images from each group. For the biggest groups of multiple thousand images, using the entire set would have resulted in each image occupying only a few pixels within the grid, leading to meaningless outputs. To be applicable to all groups we investigated grid sizes of up to 7x7 images, which is close the minimum number of images present in the smallest groups (50).
@@
In the group embedding setting, all images of a group are passed through the vision encoder of the respective model and is averaged before the caption is generated. Naturally this is only applicable to open-source models, thus not to the GPT-4V baseline.
@@
In the summary setting, we generate a caption for every individual image of the group. Then we utilize GPT-4V to summarize these captions into a single caption. We only considered unique captions. The prompt format used for this operation was:
In the following, a list of captions is provided.
Generate a caption that best describes the group of captions.
The group captions should be short and concise.
{[f"{i+1}: {caption} " for i, caption in enumerate(captions)]}
Group Caption:
@@
To ensure a fair comparison with ImageSet2Text, the grid setting necessitates a “caption curation” step. Since the models process the grid as a single image, they often generate captions containing phrases like “collage of images” or “grid of images”. These terms negatively impact performance evaluation, as the reference captions describe only the image content without mentioning grids or collages. Therefore, we removed such terms or sentences from the generated captions, evaluating only the detected content within the grid images. To not introduce any systematic bias to the evaluation, we applied this curation step to captions of all methods, also ImageSet2Text.
@@ B.3 Generating a Caption with ImageSet2Text
ImageSet2Text is not explicitly designed for group captioning, but this task serves as a key evaluation tool for assessing the accuracy of descriptions. The typical output of ImageSet2Text is a long, nuanced, and detailed textual description highlighting the main visual elements shared among the images in a given set. However, for evaluation purposes, this detailed description must be transformed into a more plain and concise textual representation, i.e., a caption.
To generate a caption from a description issued by ImageSet2Text, we utilized GPT-4V with the following prompt format:
@@
In the following, a description of a group of images is provided.
Summarize the description into a plain, single sentence caption.
Focus on the most important parts and and keep it as short as possible.
Description: {description}
Caption:
@@ B.4 Metrics
For all metrics, we used the suggested standard implementation of recent publications or if available the original authors of the papers suggesting these metrics. We use the standard rouge package to calculate the ROUGE-L (F1) score. Furthermore, we use the BLEU-4 score (sentence-wise) and the METEOR score as implemented by the nltk package. For SPICE we used the implementation of the pycocoevalcap package. For CIDEr-D, we ported the python 2 code accompanying the original publication [60] to python 3. They also provide precomputed document frequencies on the MSCOCO dataset [35], which are necessary for our task as only a single reference caption is available. For BERTScore, we utilize the bert-score package. We considered the microsoft/deberta-xlarge-mnli [23] model as basis to calculate the scores. For LLM-as-a-judge, we utilized the same GPT-40-mini model as for the rest of the evaluation with the following prompt scheme:
You are an impartial judge evaluating the equivalence of two captions.
Response only with True and False.
Caption 1: {hypothesis}
Caption 2: {reference}
Response:
Finally, for CLIPScore we utilized the Open-CLIP [27] ViT-bigG-14 model trained on the LAION2B dataset [53] (S39B B160K).
@@ B.5 Detailed Results
We provide detailed results on GroupConceptualCaptions in Tab. 6 and GroupWikiArt in Tab. 7. Additional to the metrics discussed in the main paper, we include BLEU-4 [45] for completeness. We already featured CIDEr-D and METEOR, which are designed to improve upon the weaknesses of BLEU, in our evaluation in the main paper. Furthermore, we additionally report CLIPScore for the GroupWikiArt dataset. Here, it is actually not clear if a good group captining method should attain higher scores. Depending on how the vision encoder of CLIP extracts semantic information from the image, it could be rather misleading, i.e., high scores might not indicate a good group caption. Also, if e.g., trees or bottles are visually important features, generating a description that correctly captures the style, artist and genre might not score high. Given that we construct the groups from the metadata, we do not forsee that any refernence-free metric such as CLIPScore effectively measures performance on this task.
@@
Table 6: Extended results of accuracy evaluation for in group image captioning on the GroupConceptualCaptions dataset.
@@
Model (Setting) CIDEr-D SPICE METEOR Rouge-L BLEU BERTScore LLM-Judge CLIPScore
LLaVA-1.5 (1x1 grid) 0.103 0.081 0.101 0.144 0.018 0.640 0.284 0.272
LLaVA-1.5 (2x2 grid) 0.046 0.102 0.079 0.106 0.015 0.619 0.181 0.266
LLaVA-1.5 (3x3 grid) 0.082 0.112 0.096 0.086 0.013 0.627 0.207 0.268
LLaVA-1.5 (4x4 grid) 0.092 0.121 0.110 0.089 0.014 0.632 0.198 0.273
LLaVA-1.5 (5x5 grid) 0.090 0.106 0.102 0.090 0.014 0.625 0.216 0.269
LLaVA-1.5 (6x6 grid) 0.082 0.099 0.098 0.087 0.014 0.619 0.172 0.261
LLaVA-1.5 (7x7 grid) 0.069 0.092 0.094 0.085 0.013 0.619 0.138 0.255
LLaVA-1.5 (Avg emb.) 0.053 0.041 0.074 0.107 0.014 0.586 0.103 0.232
LLaVA-1.5 (Summary) 0.038 0.085 0.112 0.091 0.012 0.626 0.198 0.301
GPT-4V (1x1 grid) 0.251 0.130 0.137 0.189 0.024 0.655 0.302 0.299
GPT-4V (2x2 grid) 0.120 0.084 0.110 0.098 0.013 0.623 0.155 0.297
GPT-4V (3x3 grid) 0.143 0.108 0.099 0.096 0.014 0.635 0.284 0.314
GPT-4V (4x4 grid) 0.146 0.105 0.104 0.099 0.013 0.649 0.276 0.315
GPT-4V (5x5 grid) 0.139 0.107 0.098 0.092 0.013 0.645 0.276 0.308
GPT-4V (6x6 grid) 0.131 0.101 0.091 0.087 0.013 0.647 0.276 0.305
GPT-4V (7x7 grid) 0.112 0.106 0.097 0.091 0.013 0.649 0.259 0.294
GPT-4V (Summary) 0.132 0.104 0.096 0.104 0.015 0.631 0.129 0.314
ImageSet2Text 0.210 0.143 0.149 0.155 0.020 0.674 0.345 0.325
Table 7: Extended results of accuracy evaluation for in group image captioning on the GroupWikiArt dataset.
@@
Model (Setting) CIDEr-D SPICE METEOR Rouge-L BLEU BERTScore LLM-Judge CLIPScore
BLIP-2 (1x1 grid) 0.002 0.034 0.055 0.063 0.011 0.539 0.019 0.255
BLIP-2 (2x2 grid) 0.046 0.060 0.060 0.091 0.017 0.542 0.038 0.263
BLIP-2 (3x3 grid) 0.117 0.103 0.077 0.088 0.015 0.583 0.179 0.303
BLIP-2 (4x4 grid) 0.092 0.102 0.070 0.076 0.014 0.577 0.132 0.295
BLIP-2 (5x5 grid) 0.068 0.091 0.065 0.058 0.010 0.565 0.132 0.291
BLIP-2 (6x6 grid) 0.052 0.079 0.047 0.047 0.009 0.551 0.085 0.278
BLIP-2 (7x7 grid) 0.062 0.084 0.046 0.049 0.009 0.551 0.057 0.273
BLIP-2 (Avg emb.) 0.004 0.054 0.076 0.086 0.013 0.576 0.028 0.294
BLIP-2 (Summary) 0.082 0.032 0.069 0.075 0.010 0.574 0.085 0.269
LLaVA-1.5 (1x1 grid) 0.001 0.012 0.033 0.043 0.006 0.542 0.000 0.219
LLaVA-1.5 (2x2 grid) 0.003 0.037 0.022 0.034 0.004 0.556 0.057 0.274
LLaVA-1.5 (3x3 grid) 0.025 0.124 0.036 0.032 0.005 0.585 0.019 0.289
LLaVA-1.5 (4x4 grid) 0.027 0.119 0.039 0.032 0.005 0.589 0.019 0.289
LLaVA-1.5 (5x5 grid) 0.028 0.115 0.036 0.031 0.005 0.583 0.028 0.293
LLaVA-1.5 (6x6 grid) 0.030 0.107 0.041 0.034 0.006 0.579 0.028 0.292
LLaVA-1.5 (7x7 grid) 0.024 0.106 0.034 0.033 0.006 0.576 0.009 0.295
LLaVA-1.5 (Avg emb.) 0.001 0.012 0.020 0.027 0.004 0.547 0.000 0.202
LLaVA-1.5 (Summary) 0.037 0.027 0.067 0.060 0.008 0.559 0.000 0.248
GPT-4V (1x1 grid) 0.002 0.012 0.064 0.064 0.008 0.558 0.028 0.242
GPT-4V (2x2 grid) 0.011 0.079 0.101 0.058 0.008 0.594 0.151 0.283
GPT-4V (3x3 grid) 0.023 0.117 0.116 0.050 0.007 0.613 0.208 0.300
GPT-4V (4x4 grid) 0.022 0.108 0.108 0.052 0.008 0.617 0.208 0.310
GPT-4V (5x5 grid) 0.021 0.140 0.108 0.053 0.009 0.621 0.160 0.301
GPT-4V (6x6 grid) 0.032 0.177 0.108 0.050 0.008 0.624 0.142 0.305
GPT-4V (7x7 grid) 0.025 0.159 0.106 0.045 0.008 0.623 0.113 0.298
GPT-4V (Summary) 0.003 0.028 0.049 0.036 0.006 0.560 0.075 0.247
ImageSet2Text 0.032 0.063 0.115 0.090 0.012 0.620 0.248 0.291
C Completeness Evaluation

### v2

@@ B.1 Dataset Composition and Construction
Given the absence of public benchmarks for evaluating ImageSet2Text in the group image captioning task, we construct two datasets. Specifically, the GroupConceptualCaptions dataset based on the ConceptualCaptions dataset (sharma2018) and the GroupWikiArt dataset based on the WikiArt dataset (artgan2018). Group sizes for these datasets are depicted in Fig. 9. The minimal size is 50, and many groups are smaller than 1000 samples, yet there are also sets in the range of 3000-4000 images. To retrieve relevant image sets from each of the two data sources, we employed different techniques.
For the ConceptualCaptions dataset (sharma2018), we grouped images that shared the same caption and applied a filter to retain only those sets with more than 100 available links in the metadata of the original dataset. This process resulted in 287 distinct captions. After downloading the images, we found that not all links were accessible. We further filtered the sets, keeping only those with at least 50 images, which resulted in 125 unique sets. We then manually reviewed all the sets, removing 9 sets that contained either duplicate images, broken images, or images that did not correspond to the caption. Additionally, we deleted any duplicate images within the remaining 116 sets. As a result, we obtained 116 image sets, with sizes ranging from 50 to 3,342 images, for a total of 23,412 images.
The initial WikiArt dataset (artgan2018) does not include explicit reference captions on which we could group the images, but it includes metadata about the artist, style, and genre. The possible values for these attributes are listed in Table 3. Using this information, we generated captions for the images based on the following rules: when aggregating by genre and style, the caption format is “<genre><genre> in <style><style> style”; when considering the artist as well, the caption format becomes “<genre><genre> by <artist><artist> in <style><style> style”. For instance, possible captions are: “Landscapes in Romanticism Style” or “Religious paintings by Edgar Degas in Impressionism Style”. To ensure sufficient data for analysis, we filtered the groups based on the number of images: groups with only two attributes (style and genre) were kept if they contained more than 499 images, while groups with three attributes (style, genre and artist) required a minimum of 50 images. After downloading the images and removing duplicates, we retained only the groups with more than 49 images. From this process, we obtained 53,707 images distributed across 105 groups. The sizes of these groups range from 50 to 4,112 images.
@@ B.2 Baseline Methods
The baselines we compare against for this task, namely BLIP-2 (li2023), LLaVA-1.5 (liu2023), GPT-4o (achiam2023) and QWEN2.5-VL (bai:2025), are not designed to process multiple images simultaneously. Therefore, we conduct experiments using different settings to enable a comparison with ImageSet2Text. These settings are referred to as (a) the grid setting, (b) the group embedding setting, and (c) the summary setting. A visual summary of these settings is provided in Fig. 10.
@@
Figure 10: Settings of utilizing VQA models to generate group captions. Light colors denote images (gray), image embeddings (purple) and output text (pink) on the level of individual images, dark colors denote aggregation on a group level. (a) is the grid setting where images are put into a collage to depict the group, (b) is the group embedding setting where embeddings of individual images after the vision encoder are averaged and (c) is the summary setting where an additional LLM is used to generate a group caption from the captions of individual instances.
@@
In the grid setting, we select images from the groups, arrange them into grids, and use these grids as inputs for the baseline models. Due to resolution constraints on the input size of the respective vision encoders, we include only a subset of images from each group. For the biggest groups of multiple thousand images, using the entire set would result in each image occupying only a few pixels within the grid, leading to meaningless outputs. To be applicable to all groups we investigated grid sizes of up to 7x7 images, which is close the minimum number of images present in the smallest groups (50).
@@
In the group embedding setting, all images of a group are passed through the vision encoder of the respective model and are averaged before the caption is generated. Naturally this is only applicable to open-source models, thus not to the GPT-4o baseline.
@@
In the summary setting, we generate a caption for every individual image of the group. Then we utilize GPT-4o to summarize these captions into a single caption. We only considered unique captions. The prompt format used for this operation is:
1In the following, a list of captions is provided.
2Generate a caption that best describes the group of captions.
3The group captions should be short and concise.
4{[f"{i+1}: {caption} " for i, caption in enumerate(captions)]}
5Group Caption:
@@
To ensure a fair comparison with ImageSet2Text, the grid setting necessitates a “caption curation” step. Since the models process the grid as a single image, they often generate captions containing phrases like “collage of images” or “grid of images”. These terms negatively impact performance evaluation, as the reference captions describe only the image content without mentioning grids or collages. Therefore, we remove such terms or sentences from the generated captions, evaluating only the detected content within the grid images. To not introduce any systematic bias to the evaluation, we applied this curation step to captions of all methods, also ImageSet2Text.
@@ B.3 Generating a Caption with ImageSet2Text
ImageSet2Text is not explicitly designed for group captioning, but this task serves as a key evaluation tool for assessing the accuracy of descriptions. The typical output of ImageSet2Text is a long, nuanced, and detailed textual description highlighting the main visual elements shared among the images in a given set. However, for evaluation purposes, this detailed description must be transformed into a more plain and concise textual representation, i.e., a caption. To generate a caption from a description issued by ImageSet2Text, we utilize GPT-4o with the following prompt format:
1In the following, a description of a group of images is provided.
2Summarize the description into a plain, single sentence caption.
3Focus on the most important parts and and keep it as short as possible.
4Description: {description}
5Caption:
@@ B.4 Metrics
For all metrics, we use the suggested standard implementation of recent publications or if available the original authors of the papers suggesting these metrics. We use the standard rouge package to calculate the ROUGE-L (F1) (lin2004) score. Furthermore, we use the BLEU-4 score (sentence-wise) (papineni2002) and the METEOR score (denkowski2014) as implemented by the nltk package. For SPICE (anderson2016) we use the implementation of the pycocoevalcap package. For CIDEr-D (vedantam2015), we ported the python 2 code accompanying the original publication (vedantam2015) to python 3. They also provide precomputed document frequencies on the MSCOCO dataset (lin2014), which are necessary for our task as only a single reference caption is available. For BERTScore (zhang2020), we utilize the bert-score package. We considered the microsoft/deberta-xlarge-mnli (he2021deberta) model as basis to calculate the scores. For LLM-as-a-judge (zheng2023), we utilized the same GPT-4o-mini model as for the rest of the evaluation with the following prompt scheme:
1You are an impartial judge evaluating the equivalence of two captions.
2Response only with True and False.
3Caption 1: {hypothesis}
4Caption 2: {reference}
5Response:
Finally, for CLIPScore (hessel2021) we utilized the Open-CLIP (ilharco2021) ViT-bigG-14 model trained on the LAION2B dataset (schuhmann2022laionb) (S39B B160K).
@@ B.5 Detailed Results
We provide detailed results on GroupConceptualCaptions in Table 4 and GroupWikiArt in Table 5, featuring both all considered metrics as well as the average rank of each method considering all different metrics. For CLIPScore on GroupWikiArt, it is actually not clear if a good group captioning method should attain higher scores. Depending on how the vision encoder of CLIP extracts semantic information from the image, it could be rather misleading, i.e., high scores might not indicate a good group caption. Also, if e.g., trees or bottles are visually important features, generating a description that correctly captures the style, artist and genre might not score high. Given that we construct the groups from the metadata, we do not foresee that any reference-free metric such as CLIPScore effectively measures performance on this task. Therefore, we excluded it when calculating the average rank on this dataset, but still report it for reference.
As discussed in the main paper, ImageSet2Text outperforms all baselines on both datasets on average.
@@
Table 4: Extended results of accuracy evaluation for in group image captioning on the GroupConceptualCaptions dataset. Average rank is calculated over the different metrics. Arrows denote if lower or higher is better.
@@
Model (Setting) CIDEr-D SPICE METEOR Rouge-L BLEU BERTScore LLM-Judge CLIPScore Average Rank
$(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\downarrow)$
LLaVA-1.5 (1x1 grid) 0.138 0.079 0.101 0.163 0.019 0.644 0.233 0.270 16.00
LLaVA-1.5 (2x2 grid) 0.145 0.115 0.101 0.150 0.020 0.657 0.276 0.271 11.31
LLaVA-1.5 (3x3 grid) 0.075 0.132 0.085 0.115 0.014 0.657 0.336 0.287 14.56
LLaVA-1.5 (4x4 grid) 0.058 0.122 0.088 0.109 0.014 0.659 0.336 0.281 15.06
LLaVA-1.5 (5x5 grid) 0.066 0.115 0.083 0.107 0.013 0.652 0.284 0.279 18.44
LLaVA-1.5 (6x6 grid) 0.066 0.121 0.084 0.099 0.013 0.651 0.302 0.274 18.06
LLaVA-1.5 (7x7 grid) 0.077 0.111 0.083 0.102 0.014 0.653 0.293 0.269 18.00
LLaVA-1.5 (Avg emb.) 0.078 0.052 0.091 0.130 0.019 0.608 0.121 0.234 21.25
LLaVA-1.5 (Summary) 0.099 0.107 0.115 0.115 0.016 0.632 0.138 0.301 17.25
Qwen2.5-VL (1x1 grid) 0.240 0.113 0.152 0.195 0.027 0.657 0.267 0.295 7.31
Qwen2.5-VL (2x2 grid) 0.168 0.118 0.140 0.153 0.017 0.651 0.216 0.309 11.00
Qwen2.5-VL (3x3 grid) 0.169 0.147 0.142 0.136 0.016 0.657 0.241 0.320 7.25
Qwen2.5-VL (4x4 grid) 0.172 0.134 0.141 0.127 0.016 0.651 0.198 0.315 10.06
Qwen2.5-VL (5x5 grid) 0.150 0.136 0.142 0.130 0.015 0.652 0.172 0.315 10.12
Qwen2.5-VL (6x6 grid) 0.121 0.135 0.146 0.130 0.015 0.657 0.250 0.313 10.38
Qwen2.5-VL (7x7 grid) 0.123 0.136 0.142 0.131 0.015 0.650 0.198 0.311 11.81
Qwen2.5-VL (Avg emb.) 0.209 0.127 0.158 0.199 0.027 0.666 0.267 0.320 4.31
Qwen2.5-VL (Summary) 0.095 0.099 0.114 0.106 0.013 0.630 0.155 0.320 19.19
GPT-4o (1x1 grid) 0.251 0.130 0.137 0.189 0.024 0.655 0.302 0.299 6.81
GPT-4o (2x2 grid) 0.120 0.084 0.110 0.098 0.013 0.623 0.155 0.297 21.31
GPT-4o (3x3 grid) 0.143 0.108 0.099 0.096 0.014 0.635 0.284 0.314 15.81
GPT-4o (4x4 grid) 0.146 0.105 0.104 0.099 0.013 0.649 0.276 0.315 14.81
GPT-4o (5x5 grid) 0.139 0.107 0.098 0.092 0.013 0.645 0.276 0.308 17.31
GPT-4o (6x6 grid) 0.131 0.101 0.091 0.087 0.013 0.647 0.276 0.305 19.56
GPT-4o (7x7 grid) 0.112 0.106 0.097 0.091 0.013 0.649 0.259 0.294 20.00
GPT-4o (Summary) 0.132 0.104 0.096 0.104 0.015 0.631 0.129 0.314 18.50
ImageSet2Text 0.210 0.143 0.149 0.155 0.020 0.674 0.345 0.325 2.50
Table 5: Extended results of accuracy evaluation for in group image captioning on the GroupWikiArt dataset. CLIPScore is added for completeness, but is not necessarily a suitable metric for this task thus kept gray. It was excluded for computing the average rank over all metrics for GroupWikiArt. Arrows denote if higher or lower is better.
@@
Model (Setting) CIDEr-D SPICE METEOR Rouge-L BLEU BERTScore LLM-Judge CLIPScore Average Rank
$(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\downarrow)$
BLIP-2 (1x1 grid) 0.002 0.034 0.055 0.063 0.011 0.539 0.019 0.255 25.29
BLIP-2 (2x2 grid) 0.046 0.060 0.060 0.091 0.017 0.542 0.038 0.263 17.21
BLIP-2 (3x3 grid) 0.117 0.103 0.077 0.088 0.015 0.583 0.179 0.303 8.36
BLIP-2 (4x4 grid) 0.092 0.102 0.070 0.076 0.014 0.577 0.132 0.295 10.64
BLIP-2 (5x5 grid) 0.068 0.091 0.065 0.058 0.010 0.565 0.132 0.291 17.07
BLIP-2 (6x6 grid) 0.052 0.079 0.047 0.047 0.009 0.551 0.085 0.278 21.79
BLIP-2 (7x7 grid) 0.062 0.084 0.046 0.049 0.009 0.551 0.057 0.273 21.57
BLIP-2 (Avg emb.) 0.004 0.054 0.076 0.086 0.013 0.576 0.028 0.294 18.21
BLIP-2 (Summary) 0.082 0.032 0.069 0.075 0.010 0.574 0.085 0.269 17.21
LLaVA-1.5 (1x1 grid) 0.001 0.001 0.033 0.059 0.010 0.538 0.000 0.220 29.79
LLaVA-1.5 (2x2 grid) 0.001 0.018 0.038 0.070 0.013 0.532 0.009 0.245 25.86
LLaVA-1.5 (3x3 grid) 0.072 0.055 0.038 0.061 0.012 0.574 0.151 0.277 18.50
LLaVA-1.5 (4x4 grid) 0.087 0.067 0.044 0.068 0.012 0.584 0.226 0.284 13.29
LLaVA-1.5 (5x5 grid) 0.092 0.070 0.051 0.070 0.013 0.589 0.226 0.287 10.57
LLaVA-1.5 (6x6 grid) 0.089 0.069 0.043 0.060 0.011 0.588 0.226 0.290 14.86
LLaVA-1.5 (7x7 grid) 0.089 0.074 0.043 0.065 0.011 0.586 0.208 0.286 13.57
LLaVA-1.5 (Avg emb.) 0.000 0.002 0.014 0.021 0.004 0.541 0.000 0.191 35.36
LLaVA-1.5 (Summary) 0.003 0.023 0.058 0.044 0.007 0.559 0.038 0.256 29.50
Qwen2.5-VL (1x1 grid) 0.001 0.031 0.074 0.061 0.007 0.576 0.057 0.252 25.00
Qwen2.5-VL (2x2 grid) 0.004 0.053 0.089 0.060 0.007 0.583 0.123 0.284 22.29
Qwen2.5-VL (3x3 grid) 0.007 0.064 0.087 0.057 0.007 0.595 0.179 0.295 19.64
Qwen2.5-VL (4x4 grid) 0.015 0.070 0.097 0.065 0.007 0.597 0.189 0.294 14.71
Qwen2.5-VL (5x5 grid) 0.014 0.068 0.103 0.065 0.008 0.599 0.274 0.304 12.86
Qwen2.5-VL (6x6 grid) 0.013 0.073 0.101 0.065 0.008 0.595 0.236 0.303 13.00
Qwen2.5-VL (7x7 grid) 0.008 0.069 0.097 0.065 0.008 0.595 0.217 0.300 15.29
Qwen2.5-VL (Avg emb.) 0.000 0.042 0.099 0.062 0.008 0.593 0.075 0.285 21.21
Qwen2.5-VL (Summary) 0.000 0.033 0.064 0.027 0.004 0.566 0.057 0.272 29.57
GPT-4o (1x1 grid) 0.002 0.012 0.064 0.064 0.008 0.558 0.028 0.242 25.50
GPT-4o (2x2 grid) 0.011 0.079 0.101 0.058 0.008 0.594 0.151 0.283 16.79
GPT-4o (3x3 grid) 0.023 0.117 0.116 0.050 0.007 0.613 0.208 0.300 13.43
GPT-4o (4x4 grid) 0.022 0.108 0.108 0.052 0.008 0.617 0.208 0.310 12.57
GPT-4o (5x5 grid) 0.021 0.140 0.108 0.053 0.009 0.621 0.160 0.301 12.00
GPT-4o (6x6 grid) 0.032 0.177 0.108 0.050 0.008 0.624 0.142 0.305 12.57
GPT-4o (7x7 grid) 0.025 0.159 0.106 0.045 0.008 0.623 0.113 0.298 14.14
GPT-4o (Summary) 0.003 0.028 0.049 0.036 0.006 0.560 0.075 0.247 29.21
ImageSet2Text 0.032 0.063 0.115 0.090 0.012 0.620 0.248 0.291 7.57
B.6 Experiments using Multi-Image Models
In addition to the main comparison to baselines as described before, we also compare against VLMs that natively allow to take multiple input images as input for producing a caption. In particular, we compare the open-source LLaVA-NeXT-Interleave (li2025llavanextinterleave) and the closed source GPT-4o-mini as used throughout this work.
We evaluated those models with a varying number of input images from the same group, which corresponded to the number of images in the grid that we used in the previous experiments. An important observation is that increasing the number of images incurs a steep increase in compute burden / API costs. For LLaVA-NeXT-Interleave, putting the full 49 images nearly consumed the whole context window of 32k, thus putting more images into the model would not be possible. For GPT-4o-mini, we have no way to check how much of the context size is used with 49 images, yet the API requests incur costs of more than 40\$ for the 221 groups we considered in our two benchmarks. Using the more capable models available through the API would approximately tenfold those costs at the current API pricing (as of October 2025).
The results are shown in Table 6 and Table 7. Again, ImageSet2Text outperforms all baselines, achieving a lower average rank across all metrics evaluated. On GroupWikiArt, LLaVA-NeXT-Interleave (25 images) approaches our performance. However, with slightly fewer or more input images, the gap widens again. In fact, we observe no consistent trend that indicates whether increasing or decreasing the number of input images benefits the baselines, suggesting that this hyperparameter is difficult to tune in practice.
Table 6: Results of accuracy evaluation for multi-image VLMs for group image captioning on the GroupConceptualCaptions dataset. Average rank is calculated over the different metrics. Arrows denote if lower or higher is better.
Model (Setting) CIDEr-D SPICE METEOR Rouge-L BLEU BERTScore LLM-Judge CLIPScore Average Rank
$(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\downarrow)$
LLaVA-NeXT (4 images) 0.214 0.103 0.113 0.165 0.023 0.665 0.405 0.285 5.25
LLaVA-NeXT (9 images) 0.141 0.102 0.100 0.148 0.019 0.668 0.362 0.296 7.25
LLaVA-NeXT (16 images) 0.129 0.109 0.102 0.143 0.018 0.671 0.379 0.300 6.38
LLaVA-NeXT (25 images) 0.121 0.104 0.100 0.149 0.016 0.671 0.345 0.294 7.56
LLaVA-NeXT (36 images) 0.053 0.101 0.119 0.136 0.013 0.642 0.328 0.296 10.12
LLaVA-NeXT (49 images) 0.077 0.100 0.120 0.149 0.016 0.636 0.293 0.296 8.50
GPT-4o (4 images) 0.155 0.111 0.120 0.135 0.018 0.650 0.250 0.309 6.88
GPT-4o (9 images) 0.145 0.111 0.117 0.117 0.015 0.657 0.233 0.318 8.75
GPT-4o (16 images) 0.147 0.109 0.120 0.117 0.015 0.657 0.190 0.319 8.25
GPT-4o (25 images) 0.159 0.116 0.120 0.121 0.015 0.656 0.207 0.320 7.19
GPT-4o (36 images) 0.159 0.117 0.117 0.123 0.015 0.660 0.207 0.321 6.31
GPT-4o (49 images) 0.159 0.115 0.122 0.121 0.016 0.652 0.198 0.321 6.75
ImageSet2Text 0.210 0.143 0.149 0.155 0.020 0.674 0.345 0.325 1.81
Table 7: Results of accuracy evaluation for multi-image VLMs for group image captioning on the GroupWikiArt dataset. CLIPScore is added for completeness, but is not necessarily a suitable metric for this task thus kept gray. It was excluded for computing the average rank over all metrics for GroupWikiArt. Arrows denote if higher or lower is better.
Model (Setting) CIDEr-D SPICE METEOR Rouge-L BLEU BERTScore LLM-Judge CLIPScore Average Rank
$(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\downarrow)$
LLaVA-NeXT (4 images) 0.004 0.012 0.039 0.067 0.013 0.552 0.009 0.230 10.14
LLaVA-NeXT (9 images) 0.022 0.054 0.064 0.065 0.013 0.587 0.104 0.268 8.14
LLaVA-NeXT (16 images) 0.063 0.069 0.078 0.099 0.017 0.606 0.104 0.283 5.14
LLaVA-NeXT (25 images) 0.069 0.069 0.087 0.097 0.017 0.609 0.113 0.292 3.86
LLaVA-NeXT (36 images) 0.026 0.067 0.049 0.059 0.007 0.561 0.123 0.288 8.00
LLaVA-NeXT (49 images) 0.022 0.058 0.050 0.057 0.007 0.548 0.104 0.275 9.43
GPT-4o (4 images) 0.006 0.064 0.098 0.055 0.010 0.610 0.236 0.287 5.86
GPT-4o (9 images) 0.010 0.065 0.085 0.039 0.007 0.606 0.208 0.285 7.86
GPT-4o (16 images) 0.012 0.080 0.089 0.043 0.007 0.604 0.170 0.284 6.43
GPT-4o (25 images) 0.006 0.076 0.085 0.040 0.006 0.603 0.179 0.283 8.21
GPT-4o (36 images) 0.019 0.082 0.087 0.038 0.006 0.608 0.179 0.290 6.50
GPT-4o (49 images) 0.007 0.076 0.087 0.036 0.005 0.604 0.217 0.285 8.00
ImageSet2Text 0.032 0.063 0.115 0.090 0.012 0.620 0.248 0.291 3.43
B.7 Compute Infrastructure
The experiments investigated in this section were performed on a mix of A40 (48GB) and A100 (80GB) cards, depending on availability on the cluster. Individual nodes featuring 8 GPUs were equipped with 1-2TB of RAM and AMD EPYC CPUs with 64/128 total CPU cores. Calls to OpenAI GPT models were performed using the latest API version as of March 2025. We provide the full code along with all required dependencies. However, the key package versions are as follows:
  • python = 3.10
  • pytorch = 2.6.0+cu124
  • transformers = 4.49
Appendix C Completeness Evaluation